What is 7n-4n+8=2n+3n

Answer by 100601(6)

7n-4n+8=2n+3n
Do 7n-4n and 2n+3n to simplify the equation:
3n+8=5n
Subtract 3n on both sides:
3n-3n+8=5n-3n
8=2n
Divide 8 by 2 to isolate the n variable:
n=4